# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

### v1

@@ 4.2 Outcomes Decomposition Rule
Under the assumptions of Theorem 4.4, we have
@@
$+\sqrt{\frac{\sigma^{2}+\sigma_{t}^{2}}{{\left|{\mathcal{S}}\right|}}}{\left\|g_{t}(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X})\right\|}J_{t}+\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\epsilon^{i}_{t}.$
@@
While Theorem 4.2 establishes that distinct subpopulations yield different state evolution observations, Corollary 4.7 reveals two crucial implications for finite-sample analysis. First, the outcome distribution within each subpopulation incorporates unit-level heterogeneities, as demonstrated by the first two terms in Eq.(11). Second, the outcome uncertainty terms—represented by the final two expressions in Eq.(11)—decay at a rate of $\sqrt{\left|{\mathcal{S}}\right|}$. The subsequent sections examine these implications in detail, showing how DPNB addresses such observational complexities while generating additional samples.
4.3 Unit-level Heterogeneities
Experimental units exhibit distinct covariates, such as age and gender. Network interference also emerges through diverse local patterns, jointly inducing heterogeneous characteristics. These heterogeneities are captured in (11) through variations in $\mu^{ij}$, $\mu_{t}^{ij}$, and ${\vec{X}}^{i}$. This foundation motivates our strategy of selecting members for each subpopulation $\mathcal{S}$ solely based on treatment allocation. Specifically, this approach enables us to treat $\mathcal{S}$ as a random sample, supporting the assumption that each subpopulation represents the original experimental population. Indeed, Theorem 4.2 demonstrates the effectiveness of this batching strategy, evidenced by the invariance of $\varphi_{t}$ across subpopulations. Meanwhile, the result of Corollary 4.7 suggests maximizing subpopulation sizes to leverage stronger smoothing effects through averaging over larger groups of units.
4.4 Endogenous Observation Noise
The noise terms in the outcome decomposition rule exhibit complex correlation structures. According to Theorem 4.4, the random vectors $\vec{R}_{0},\ldots,\vec{R}_{T-1}$ in Eq. (10) asymptotically follow a centered Gaussian distribution as population size $N$ increases, thus functioning as noise terms. These vectors, however, demonstrate a complex endogenous dependency structure through several mechanisms.
Initially, for each $t$, the random vector $\vec{R}_{t}$ derives its randomness from $\vec{Z}_{0},\vec{Z}_{1},\ldots,\vec{Z}_{t}$. Consequently, the random vectors $\vec{R}_{0},\ldots,\vec{R}_{T-1}$ exhibit correlation, generating temporal correlation in the observational noise.
Additionally, in finite populations where $N$ is bounded, the elements of $R^{i}_{t}$ demonstrate cross-unit correlation. The magnitude of this correlation is governed by the constant $c$ in Theorem 4.4, which varies across settings and remains unidentified.
Furthermore, the potential outcome of unit $i$ at time $t$ denoted by $Y^{i}_{t}{}(\bm{W}_{t}{})$ incorporates $R^{i}_{t-1}$, which correlates with elements of $\vec{R}_{t}$; this introduces correlation between noise terms and unit outcomes. The strength of this correlation depends on the magnitudes of $\sigma$ and $\sigma_{t}$, which remain unknown and can vary substantially across different settings.
These three factors collectively define a complex endogenous dependency structure, indicating how uncertainty in the interference structure can compromise outcome measurements and introduce substantial bias when overlooked. The following example examines an autoregressive model with unanticipated unit interactions, demonstrating how even minimal second-order interference induces endogeneity and generates biased estimates.
Example 4.8
Consider a simplified version of the outcome specification in Eq. (3) where $g_{t}(Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1}{},{\vec{X}}^{i})=Y^{i}_{t}{}(\bm{W}_{t}{})$ and $h_{t}{}(Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1}{},{\vec{X}}^{i})=\Theta^{i}Y^{i}_{t}{}(\bm{W}_{t}{})$, with $\mu^{ij}+\mu_{t}^{ij}=0$ for all $i$, $j$, and $t$. These conditions reflect the available partial information confirming negligible interference between units. Also, assume $\epsilon^{i}_{t}$ is independent from all other variables. We can write:
$Y^{i}_{t+1}{}(\bm{w}_{t+1}{})=\Theta^{i}Y^{i}_{t}{}(\bm{w}_{t}{})+\sqrt{\sigma^{2}+\sigma_{t}^{2}}{\left\|\vec{Y}_{t}{}(\bm{w}_{t}{})\right\|}R^{i}_{t}+\epsilon^{i}_{t}.$
To estimate $\Theta^{i}$ for a specific unit $i$, we employ ordinary least squares regression, regressing outcomes $Y^{i}_{1}{}(\bm{w}_{1}{}),\ldots,Y^{i}_{T}{}(\bm{w}_{T}{})$ on their corresponding lagged values $Y^{i}_{0}{}(\bm{w}_{0}{}),\ldots,Y^{i}_{T-1}{}(\bm{w}_{T-1}{})$. Under the asymptotic regime where $N\rightarrow\infty$, we characterize the bias of estimator $\hat{\Theta}^{i}$ as follows:
$\mathbb{E}\left[\hat{\Theta}^{i}\Big{|}\bm{Y}{}(\bm{w}{})\right]-\Theta^{i}$ $=\frac{1}{\sum_{t=0}^{T-1}Y^{i}_{t}{}(\bm{w}_{t}{})^{2}}\sum_{t=0}^{T-1}\sqrt{\sigma^{2}+\sigma_{t}^{2}}{\left\|\vec{Y}_{t}{}(\bm{w}_{t}{})\right\|}Y^{i}_{t}{}(\bm{w}_{t}{})\mathbb{E}\left[R^{i}_{t}\big{|}\bm{Y}{}(\bm{w}{})\right].$
Based on the results of Corollary 4.7, the magnitude of this complex noise diminishes as the subpopulation size $\left|{\mathcal{S}}\right|$ increases, supporting the selection of larger subpopulations. Therefore, in the extreme case, when $\mathcal{S}$ encompasses the entire population ($\left|{\mathcal{S}}\right|=N$), the noise magnitude decays at a rate of $\sqrt{N}$.
4.5 How to use DPNB to generate efficient samples?
Both unit-level heterogeneities and endogenous observation noise suggest maximizing subpopulation sizes, which would ideally lead to selecting the entire experimental population. This approach aligns with the proposal in Section 3.4, where we observe a single instance of state evolution. Therefore, this presents a nuanced trade-off: larger subpopulations increase overlap between groups, reducing variation across subpopulations and thereby decreasing the number of effective samples. Given that the magnitude of endogenous noise varies across different settings, we aim to develop a data-driven methodology to address this trade-off optimally. This leads us to propose a cross-validation method, which we examine in detail in the subsequent section.
@@ 5 Counterfactual Cross Validation
The precision of estimated counterfactuals, as established in Theorem 3.6, depends on accurately estimating the state evolution mappings $f_{t}$. This estimation process faces two main challenges that require careful validation:
@@
The true specification of potential outcomes is often unknown, requiring us to approximate the state evolution mappings $f_{t}$. This approximation involves two related aspects. First, we must choose appropriate summary statistics of the joint distribution of outcomes and treatments to serve as inputs to our model - an approach analogous to feature engineering in supervised learning. For example, Shirani and Bayati, (2024) used sample means of outcomes, treatments, and their products. While domain knowledge can guide this selection, we need a systematic way to validate these choices. Second, depending on the complexity of experimental settings, we may need to employ a range of estimation techniques, from simple linear regression to more sophisticated methods such as neural networks. The choice of technique and its specific implementation (e.g., architecture, hyperparameters) must be validated to prevent issues like model instability and estimation bias. These two aspects are linked, as both contribute to how well we can approximate the true mapping $f_{t}$. A data-driven validation methodology helps us jointly optimize these choices while reducing misspecification risks.
@@
As established in §4.5, DPNB offers an effective strategy for addressing unit-level heterogeneities and controlling endogenous noise through batching. However, the choice of batch size presents a bias-variance trade-off. Larger batches better average out heterogeneities and noise, reducing estimation bias, but they also decrease the number of distinct batches available for learning, increasing estimation variance. Conversely, smaller batches provide more samples for learning, reducing variance, but may retain more heterogeneity and noise, potentially introducing bias. Beyond batch size, the number of batches creates a similar trade-off that interacts with the batch size selection. Therefore, data-driven validation helps find the optimal combination of batch size and number of batches that balances these competing effects in a given experimental context.
@@ 5.1 Counterfactual Cross Validation Algorithm
The parameters being selected through this cross-validation include both the choice of estimator from a candidate set $\mathfrak{L}_{\mathsf{E}}$ and the batching configuration from a candidate set $\mathfrak{L}_{(\mathsf{s_{b}},\mathsf{n_{b}})}$. Each estimator $\mathsf{E}\in\mathfrak{L}_{\mathsf{E}}$ represents a specific combination of feature generation (what summary statistics to use) and estimation method (e.g., linear regression or neural network). The batching parameters $(\mathsf{s_{b}},\mathsf{n_{b}})\in\mathfrak{L}_{(\mathsf{s_{b}},\mathsf{n_{b}})}$ specify the size $\mathsf{s_{b}}$ and number $\mathsf{n_{b}}$ of training batches. For each held-out time block, we train models using all possible combinations of these parameters on the remaining blocks and evaluate their performance on the validation batches obtained from the held-out block.
Step 1: Reference counterfactual construction
After the experiment is completed, we obtain the following data. The treatment matrix $\bm{w}{}$ and observed outcomes matrix, $\bm{Y}{}(\bm{w}{})$. We first compute the sample mean of outcomes over time as ground truth (counterfactual) for evaluating estimations in subsequent steps. More precisely, for each time period $t\in\{0,1,\ldots,T\}$, and each validation batch $\mathcal{S}_{j}^{v}$ with $j\in\{1,\ldots,b_{v}\}$, we calculate average of outcomes at time $t$ across units in $\mathcal{S}_{j}^{v}$ and denote that by $\text{CFE}_{t}^{\mathcal{S}_{j}^{v}}$.
@@
Treatment allocation $\bm{w}{}$, observed outcomes $\bm{Y}{}(\bm{w}{})$, validation batches $\{\mathcal{S}_{j}^{v}\}_{j=1}^{b_{v}}$, time blocks $\mathfrak{L}_{\mathsf{tb}}$, loss function $\ell:\mathbb{R}^{(T+1)b_{v}}\times\mathbb{R}^{(T+1)\times b_{v}}\to\mathbb{R}_{+}$, and candidate estimators $\mathfrak{L}_{\mathsf{E}}$ and batch parameters $\mathfrak{L}_{(\mathsf{s_{b}},\mathsf{n_{b}})}$
Step 1: Reference Counterfactual Construction
$\big{\{}\text{CFE}_{t}^{\mathcal{S}_{j}^{v}}\big{\}}_{t=0}^{T}\leftarrow$ mean of $\bm{Y}{}(\bm{w}{})$ for units belonging to $\mathcal{S}_{j}^{v},\;j=1,\ldots,b_{v}$
@@
Algorithms 5 and 7, presented in Appendices 10 and 11 respectively, provide examples of candidate estimators based on linear regression.
Remark 5.2
In experimental settings with strong temporal patterns such as seasonality, we first detrend the observed outcomes, e.g., see Algorithm 7 in Appendix 11. This procedure augments Step 2 of Algorithm 2 as follows. For each estimator $\mathsf{E}$, we utilize the complete observed dataset to estimate a baseline counterfactual—defined as the counterfactual outcome for all units under control—using an estimation model optimized for temporal pattern detection. We then generate filtered data by subtracting this baseline counterfactual from observed outcomes. This filtered data serves as input for the main estimator, which focuses specifically on treatment effect identification. The final estimates are obtained by combining the baseline counterfactual with the estimated treatment effects. The consistency of this estimation approach relies on a structural assumption of weak additive separability between baseline outcomes and treatment effects.
6 Benchmark Toolbox
@@
We evaluate our framework using six semi-synthetic experiments that combine simulated environments with real-world data. This approach offers two key advantages: it maintains realistic data characteristics while allowing us to compute ground truth values for our estimands. Unlike real experimental settings where outcomes are observed under a single scenario, these settings provide ground truth values for any desired scenario. This enables rigorous evaluation of estimation methods under realistic conditions. In the following sections, we specify the treatment allocation, outcomes, and network structure for each experimental setting.
6.1 LLM-based Social Network

### v2

@@ 4.3 Navigating Estimation Trade-offs with DPNB
Under the assumptions of Theorem 4.4, for any subpopulation $\mathcal{S}$, we have
@@
$+\sqrt{\frac{\sigma^{2}+\sigma_{t}^{2}}{{\left|{\mathcal{S}}\right|}}}{\left\|g_{t}(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X})\right\|}J_{t}+\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\epsilon^{i}_{t},$
@@
This corollary mathematically grounds a trade-off between two estimation extremes. The third term in the right-hand side of Eq. (16) shows that the standard deviation of the endogenous network heterogeneity noise scales with $1/\sqrt{\left|{\mathcal{S}}\right|}$. At one extreme, a unit-level estimation (ULE) uses each unit as a subpopulation of size one ($\left|{\mathcal{S}}\right|=1$). This maximizes the number of observations ($N\times T$) but makes the model susceptible to the large endogenous noise term, which can bias estimation. At the other extreme, a population-level estimation (PLE) uses the entire population as a single batch ($\left|{\mathcal{S}}\right|=N$). This maximally suppresses the noise but leaves only a single time series ($T$ data points), limiting statistical power and increasing the risk of misspecification.
Our proposed distribution-preserving network bootstrap method provides a flexible framework that interpolates between these two extremes. By generating multiple samples from subpopulation averages, it aims to increase the effective sample size beyond $T$ while still leveraging the noise-reducing benefits of outcome aggregations.
However, the choice of batch configuration itself introduces a nuanced bias-variance trade-off. Larger batches behave more like PLE. They more effectively average out the endogenous noise, reducing estimation bias. But this comes at the cost of generating fewer, less variable batches, which can limit statistical power and make it difficult to identify complex functional forms. Smaller batches provide more samples and greater cross-sectional variation, which is helpful for learning the underlying dynamics and enhancing statistical power. But, these smaller averages are more susceptible to the endogenous noise, potentially reintroducing the bias that impacts unit-level models.
A formal characterization of the optimal batching strategy is a significant challenge, as it depends on features of the experimental environment that are typically unknown, such as the relative magnitudes of mean interference ($\mu$) and network heterogeneity ($\sigma$). While rigorous theoretical analysis of these trade-offs represents a promising direction for future research, we take a data-driven approach in the next section. We embrace the complexity of this trade-off and use DPNB to generate multiple observational trajectories from a single experiment, enabling our cross-validation framework with more samples of the state evolution equation. This framework forms the core of our proposed methodology for settings with unknown networks. Before detailing this framework, we briefly discuss how our approach can be adapted when partial network information is available.
4.4 Incorporating Partial Network Knowledge
Settings where the network structure is entirely unknown represent a challenging scenario common in many experiments. However, when partial knowledge of the network is available, this information should be leveraged to improve estimation accuracy. Eq. (12) provides a decomposition of unit outcomes where the first term captures the interference effect. We can adjust this term to reflect any available knowledge about the interference network, as illustrated in the following example.
Example 4.6 (Known clusters)
Consider a scenario where the network structure is partially known with clusters $C^{1},\ldots,C^{K}$ (e.g., the experimental population consists of individuals from different villages or online communities). In this context, the interference term in Eq. (12) can be rewritten to distinguish between within-cluster and between-cluster effects:
$Y^{i}_{t+1}{}(\bm{W}_{t+1}{})=$ $\frac{1}{N}\sum_{k=1}^{K}\Bigg[\underbrace{\sum_{j\in C^{k},i\notin C^{k}}\left(\mu^{ij}+\mu_{t}^{ij}\right)g_{t}\left(Y^{j}_{t}{}(\bm{W}_{t}{}),W^{j}_{t+1}{},{\vec{X}}^{j}\right)}_{\text{Between-cluster interference}}$
$+\underbrace{\sum_{i,j\in C^{k}}\left(\mu^{ij}+\mu_{t}^{ij}\right)g_{t}\left(Y^{j}_{t}{}(\bm{W}_{t}{}),W^{j}_{t+1}{},{\vec{X}}^{j}\right)}_{\text{Within-cluster interference}}\Bigg]$
where we would typically expect the magnitude of $\left(\mu^{ij}+\mu_{t}^{ij}\right)$ to be larger when units $i$ and $j$ belong to the same cluster, reflecting stronger intra-cluster ties.
The reformulated decomposition in the example above has practical implications for estimation. Instead of relying on a single population-level average (as in PLE) or averages over arbitrary subpopulations (as in DPNB), one can now leverage the known cluster structure. By computing outcome and treatment averages for each cluster, we can generate a richer set of summary statistics to learn the state evolution dynamics. Ultimately, whether operating with no network knowledge or incorporating partial structural information, our proposed methodology provides the necessary data for robust estimation. The following sections detail the cross-validation framework that leverages this data and then demonstrate the performance of our framework empirically.
@@ 5 Counterfactual Cross Validation
The precision of estimated counterfactuals, as established in Theorem 3.5, depends on accurately estimating the SE mappings $f_{t}$. This estimation process faces two main challenges that require systematic validation:
@@
The true specification of potential outcomes is often unknown, requiring us to approximate the SE mappings $f_{t}$. This approximation involves two related aspects. First, we must choose appropriate summary statistics of the joint distribution of outcomes and treatments to serve as inputs to our model, an approach analogous to feature engineering in supervised learning. For example, Shirani and Bayati, (2024) used sample means of outcomes, treatments, and their products. While domain knowledge can guide this selection, we need a systematic way to validate these choices. Second, depending on the complexity of experimental settings, we may need to employ a range of estimation techniques, from simple linear regression to more sophisticated methods such as neural networks. The choice of technique and its specific implementation (e.g., architecture, hyperparameters) must be validated to prevent issues like model instability and estimation bias. These two aspects are interconnected, as both contribute to how well we can approximate the true SE mappings $f_{t}$. A data-driven validation methodology helps us jointly optimize these choices while reducing misspecification risks.
@@
As established in §4.3, DPNB offers an effective strategy to control endogenous noise by batching experimental units. However, the choice of batch size presents a trade-off. Larger batches better average out noise, but decrease the number of distinct batches available for learning. Conversely, smaller batches provide more samples for learning, but may retain more noise that potentially introduces bias. The number of batches creates a similar trade-off that interacts with batch size selection. Therefore, data-driven validation is essential to find the optimal combination of batch size and number of batches that balances these competing effects within a given experimental context.
@@ 5.1 Counterfactual Cross Validation Algorithm
The parameters being selected through the cross-validation include both the choice of estimator from a candidate set $\mathfrak{L}_{\mathsf{E}}$ and the batching configuration from a candidate set $\mathfrak{L}_{(\mathsf{s_{b}},\mathsf{n_{b}})}$. Each estimator $\mathsf{E}\in\mathfrak{L}_{\mathsf{E}}$ represents a specific combination of feature generation (what summary statistics to use) and estimation method (e.g., linear regression or neural network). The batching parameters $(\mathsf{s_{b}},\mathsf{n_{b}})\in\mathfrak{L}_{(\mathsf{s_{b}},\mathsf{n_{b}})}$ specify the size $\mathsf{s_{b}}$ and number $\mathsf{n_{b}}$ of training batches (e.g., 100 batches each containing 20% of units or 500 batches each containing 10% of units). For each held-out time block, we train models using all possible combinations of these parameters on the remaining blocks and evaluate their performance on the validation batches obtained from the held-out block.
Figure 5: Counterfactual Cross-Validation. Time horizon is partitioned into blocks for leave-one-out validation. Models are trained on remaining blocks and evaluated via MSE to select optimal configurations.
Step 1: Reference ground truth construction
After the experiment is completed, we obtain the treatment matrix $\bm{w}{}$ and observed outcomes panel $\bm{Y}{}(\bm{w}{})$. We first compute the sample mean of outcomes over time as ground truth for evaluating estimations in subsequent steps. More precisely, for each time period $t\in\{0,1,\ldots,T\}$, and each validation batch $\mathcal{S}_{j}^{v}$ with $j\in\{1,\ldots,b_{v}\}$, we calculate average of outcomes at time $t$ across units in $\mathcal{S}_{j}^{v}$ and denote that by $\text{CFE}_{t}^{\mathcal{S}_{j}^{v}}$.
@@
Algorithm 2 presents our counterfactual cross-validation procedure. To ensure consistency of the final selected estimator, we construct the initial estimator list $\mathfrak{L}_{\mathsf{E}}$ such that each candidate estimator is consistent under certain assumptions. Our cross-validation algorithm then optimizes estimator selection for real-world scenarios with finite experimental units. For example, we can use variations of Algorithm 1 based on Ridge regression with different penalty terms as candidate estimators, with our cross-validation optimizing the Ridge penalty selection. Similarly, Algorithm 6 in Appendix 11 provides additional candidate estimators based on linear regression that extend Algorithm 1 by incorporating subpopulation-specific terms. In this case, the optimization process also involves selecting the optimal batch configuration.
When additional data are available (e.g., multiple clusters or pre-experiment observations), Algorithm 2 can be enhanced to leverage these datasets for more sophisticated training and validation.
@@
Treatment allocation $\bm{w}{}$, observed outcomes $\bm{Y}{}(\bm{w}{})$, validation batches $\{\mathcal{S}_{j}^{v}\}_{j=1}^{b_{v}}$, time blocks $\mathfrak{L}_{\mathsf{tb}}$, loss function $\ell:\mathbb{R}^{(T+1)\times b_{v}}\times\mathbb{R}^{(T+1)\times b_{v}}\to\mathbb{R}_{+}$, and candidate estimators $\mathfrak{L}_{\mathsf{E}}$ and batch parameters $\mathfrak{L}_{(\mathsf{s_{b}},\mathsf{n_{b}})}$
Step 1: Reference Ground Truth Construction
$\big\{\text{CFE}_{t}^{\mathcal{S}_{j}^{v}}\big\}_{t=0}^{T}\leftarrow$ average of observed outcomes over time for units belonging to $\mathcal{S}_{j}^{v},\;j=1,\ldots,b_{v}$
@@
In experimental settings with strong temporal patterns such as seasonality, we first detrend the observed outcomes, following Assumption 3.5. We can incorporate detrending into Algorithm 2 by augmenting Step 1 as follows:
  • Fit an auxiliary model on the full observed data to capture temporal patterns and compute the baseline counterfactual as the sample mean of outcomes under all-control.
  • Create filtered data by subtracting the baseline counterfactual from observed outcomes.
  • Feed the filtered data into Step 2 that targets treatment-effect identification.
  • Combine the baseline counterfactual with estimated treatment effects for final estimates.
Algorithm 8 in Appendix 12 presents an example of counterfactual estimation with detrending, where the consistency result relies on a weak additive separability assumption between baseline dynamics and treatment effects (see Appendix 12).
6 Interference Gym
@@
In this section, we introduce six semi-synthetic experiments that combine simulated environments with real-world data (Table 1). This approach offers two key advantages: it maintains realistic data characteristics while allowing us to compute ground truth values for our estimands. Unlike real experimental settings where outcomes are observed under a single treatment allocation, these settings provide ground truth values for any desired scenario. This enables rigorous evaluation of estimation methods under realistic conditions. We specify the treatment allocation, outcomes, and network structure for each experimental setting below. Each setting is accompanied by a corresponding Jupyter notebook accessible via these icons: [Uncaptioned image] . These notebooks are part of our Python package and allow running our estimation framework or extracting data for other applications.
LLM-based Social Network – Social media platform (like Facebook) with AI agents as users
Treatment Feed ranking: random ordering (control) vs. friend engagement weighting (treatment)
Outcome User engagement metrics (sum of likes and replies per user)
Interference Structure Preferential attachment model with directed follower-following relationships
Key Characteristic AI agents with demographically diverse personas based on US Census data
Belief Adoption Model – Opinion diffusion through social coordination games
Treatment Promotional campaign to promote Opinion A (new belief)
Outcome Binary adoption decision (Opinion A vs. Opinion B)
Interference Structure Coordination game with payoff-based decisions influenced by friends’ opinions
Key Characteristic Pokec social network (3 regional networks: 3,366-42,971 users) with demographic profiles
NYC Taxi Routes – Urban transportation network with route dependencies
Treatment Pricing experimentation on randomly selected routes
Outcome Trip count per route during 6-hour periods
Interference Structure Route adjacency based on geographic proximity and functional relationships
Key Characteristic Real temporal pattern using NYC taxi records (58M trips, Jan-Mar 2024)
Exercise Encouragement – Digital health intervention with social influence
Treatment Digital motivational messages for physical activity
Outcome Binary exercise decision (Bernoulli distribution)
Interference Structure Twitter social circles with peer influence (avg. 21.74 connections)
Key Characteristic Demographic-based response patterns with weekly cycles based on US Census data
Data Center – Server farm with workload redistribution effects
Treatment Processing power enhancements for selected servers
Outcome Server utilization rate (proportion of time server remains busy during each interval)
Interference Structure Join-the-shortest-queue routing creates implicit interference patterns
Key Characteristic Time-dependent Poisson arrivals with realistic daily pattern and job classification
Auction Model – Competitive auction market with bidders competing for different objects
Treatment Promotional interventions increasing bidder valuations by $\tau$% for selected objects
Outcome Value of each object in each auction round
Interference Structure Indirect effects through strategic bidding and price-driven market dynamics
Key Characteristic Market-based pricing with heterogeneous bidder types and realistic valuation strategies
Table 1: Description of Experimental Environments
6.1 LLM-based Social Network [Uncaptioned image]